Grade 12th passMechanics

A particle Travels half the distance of a straight German with a speed V metre per second the remaining part of the distance is covered with speed 4 6 m per second other half of the remaining the average speed of the particle is

Zsinab
8 Years agoGrade 12th pass

1 Answer

Deepti
8 Years ago
let both the halfs of distances be D. speed=distance/time
average distance= D + D / time for 1st half + time for 2nd half ….. (i)
(time for 1st half = D/v , time for 2nd half = D/4.6)
so the total time becomes 4.6D+vD/4.6v
subsituting the value of total time and solving we get average speed = 2*4.6V/4.6+v
(the D gets cancelled)
Hope this helped.